# Quantum Networks: from Multipartite Entanglement to Hypergraph Immersion

### v1

@@ Abstract
Multipartite entanglement, a higher-order interaction unique to quantum information, offers various advantages over bipartite entanglement in quantum network (QN) applications. Establishing multipartite entanglement across remote parties in QN requires entanglement routing, which irreversibly transforms the QN topology at the cost of existing entanglement links. Here, we address the question of whether a QN can be topologically transformed into another via entanglement routing. Our key result is an exact mapping from multipartite entanglement routing to Nash-Williams’s graph immersion problem, extended to hypergraphs. This generalized hypergraph immersion problem introduces a partial order between QN topologies, permitting certain topological transformations while precluding others, offering discerning insights into the design and manipulation of higher-order network topologies in QNs.
@@
Quantum entanglement, a pivotal quantum resource [1], allows for the simultaneous correlation of multiple qubits. Multipartite entanglement offers notable benefits, in terms of both efficiency and scalability, over bipartite entanglement in a variety of quantum information applications such as sharing secret among multiple parties [2] and reducing the memory requirements for entangling qubits [3]. These benefits make multipartite entanglement a necessary building block for future quantum networks (QN) [4, 5, *multipartite-q-netw_ctpv21, 7]. In a QN, each vertex (node) typically represents a local assembly of qubits entangled with qubits from other vertices, and each edge (link) signifies an entangled state between qubits that belong to different vertices [8, *conpt_mgh21]. In this setting, multipartite entanglement is most effectively represented as a higher-order interaction, or in the language of hypergraph theory, as a hyperedge [10]. Contrasting with ordinary edges that connect just two vertices, hyperedges can connect several vertices at once—a feature that has wider applications in not only physics but also chemistry [11], biology [12], and social sciences [13, *krumov2011collab, *bianconi2021hyper, *Kwang2023core].
@@
In ordinary QNs (without hyperedges), vertices not directly connected by an edge can still be entangled through a fundamental operation known as entanglement routing [17, *q-netw-route_pkttjbeg19, *q-netw-route_lphnm20, 20], provided that there is a path of edges connecting the vertices. This operation, however, utilizes all the edges along the path, subsequently deleting them from the QN, leading to a dynamic process called path percolation [21]. As a result, each entanglement routing operation uniquely modifies the QN’s topology in an irreversible manner, permitting certain types of topological transformations while precluding others, raising the question: can a QN’s topology be transformed from one to another through a series of entanglement routing operations? It is also unknown how this question extends to hypergraph QNs due to the ambiguous definition of paths involving multipartite entanglement [20]. Understanding this question is key to grasping the complexity of many-body entanglement structures (e.g., matrix product states [22], which are derivable from QNs [7, 23]). Moreover, the insights can be applied to evaluate the effectiveness of qubit layouts (e.g., the hex lattices used by IBM [24]). This is especially relevant as small-scale multipartite-entanglement QNs have been implemented on superconducting [25, *chen2014qubit] and optical [27, *kok2007linear] platforms.
@@
Here, we address this question of entanglement routing, involving both edges and hyperedges, by establishing an exact mapping to a topological problem on hypergraphs. We first identify and simplify entanglement routing into two hypergraph rules by extending the routing operation to multipartite entanglement. Then, we show that the corresponding topological problem can be viewed as a generalization of the well-established graph immersion problem [29], which, to our best knowledge, has not been generalized to hypergraphs. This aptly allows us to formalize our problem as hypergraph immersion, which establishes a partial order [30] between hypergraph topologies. Following graph immersion [31], we also prove in an accompanying work that the algorithmic complexity of solving hypergraph immersion is polynomial  [^1] In preparation.. This unique bridge between quantum information and graph theory may provide some useful insights into the efficient design of multipartite QNs.
@@
Figure 1: Hypergraph immersion. (a) The “coalescence” operation merges two hyperedges that share at least one vertex, resulting in a new hyperedge that joins all vertices from the original two. (b) The “evaporation” operation disconnects a vertex from a hyperedge, reducing the hyperedge’s size by one. (c) Transformation of QN from an initial topology to a final one, achieved through a series of operations. The final topology is effectively immersed within the original.
@@
These states, albeit varying in size ($r\geq 2$), always have the lowest nontrivial generalized Schmidt rank [33, *eisert2001schmidt] to exhibit entanglement ($\text{rank}=2$). GHZ states are prevalent in quantum systems, such as the random transverse-field quantum Ising model, where the critical ground state consists exclusively of GHZ-state magnetic domains [35]. The Ising configuration naturally facilitates a QN interpretation using complex network theory [36].
How do we entangle multiple vertices by routing via several hyperedges? Given, for example, two GHZ states of $r_{1},r_{2}$ qubits respectively and at least one common vertex that shares two qubits separately from the two states, a simple projector on the two qubits,
@@
on the vertex, yielding a smaller GHZ state with $r-1$ qubits. With appropriate rotations of |$\ket{0}$ and |$\ket{1}$, these two operations are also extensible to general Schmidt-rank-2 states of the form $|\alpha_{1}\otimes\alpha_{2}\otimes...\otimes\alpha_{n}\rangle+z|\beta_{l}\otimes\beta_{2}\otimes...\otimes\beta_{n}\rangle$ [37]. An example of such is the cat state, commonly used in continuous-variable optical communications [38].
@@
Coalescence and evaporation operations.—The topological emphasis of Eqs. (2) and (3) suggests that they can be simplified as hypergraph operations. Consider two hypergraphs $H$ and $G$, each comprising sets of vertices $V(H)$ and $V(G)$, and sets of hyperedges $E(H)$ and $E(G)$, respectively. Both $H$ and $G$ are loopless and undirected, but may include multi-hyperedges. The feasibility of deriving $H$ from $G$ through entanglement routing is equivalent to asserting that $H$ can be derived from $G$ through the following hypergraph operations:
@@
  • Evaporation: disconnecting an hyperedge from one vertex that it is incident on.
One can visualize vertices in the hypergraph as spatially distributed objects (pink) that are “hydrophilic” (Fig. 1). Correspondingly, each hyperedge $e$ can be visualized as a water droplet, touching those hydrophilic objects (vertices) that are incident with $e$. In this physical analogy, the coalescence operation results in the merging of two droplets that share contact with a common vertex $v$, forming a larger droplet [Fig. 1]; the evaporation operation simulates the reduction in volume of a droplet and the loss of its connections to the vertices it was previously in contact with [Fig. 1]. This physical analogy motivates our naming of the two operations.
Note that both operations are irreversible: the coalescence reduces the number of hyperedges by one; the evaporation reduces the number of vertices incident with the hyperedge by one. Hence, for $H$ to be derived from $G$ through these operations, a necessary condition is $\left|V(H)\right|\leq\left|V(G)\right|$ and $\left|E(H)\right|\leq\left|E(G)\right|$, i.e., $H$ must not possess more vertices or hyperedges than $G$. In general, determining whether $H$ can be derived from $G$ poses a nontrivial problem. In the following, we present an alternative definition in parallel to these operations.
@@
In other words, $\alpha$ is an injective mapping from vertices in $H$ to vertices in $G$, and from hyperedges in $H$ to edge-disjoint connected subgraphs in $G$. The formulation of $\alpha$ mirrors the graph immersion as outlined in Ref. [29], which has a similar definition that maps edges in $H$ to edge-disjoint paths in $G$, but with $H$ and $G$ being ordinary graphs. For ordinary edges, the function $\alpha$ reduces to the definition of graph immersion.
@@
An immersion $\alpha$ of hypergraph $H$ in hypergraph $G$ exists if and only if $H$ can be obtained from a subgraph of $G$ by a sequence of coalescence and evaporation operations.
@@
If an immersion $\alpha$ exists, we can show that $H$ is isomorphic to a graph $H^{\prime}$ obtained by applying the coalescence and evaporation operations on a subgraph of $G$, where the bijection corresponding to the isomorphism $f:V(H)\to V(H^{\prime})$ takes the same value as $\alpha$, i.e., $\forall v\in V(H)$, $f(v)=\alpha(v)$.
We obtain $H^{\prime}$ as follows. We focus on the subgraph of $G$, $G^{\prime}\subseteq G$ s.t. $G^{\prime}=\bigcup_{e\in E(H)}\alpha(e)$. By the definition of immersion $\alpha$, $G^{\prime}$ is necessarily connected. For each $e\in E(H)$ with distinct ends $v_{1},v_{2},\dots,v_{\absolutevalue{e}}$, we then apply the coalescence operation to pairs of (hyper)edges in $\alpha(e)\subseteq G$ until we obtain one single hyperedge with all vertices in $\{\alpha(v_{1}),\alpha(v_{2}),\dots,\alpha(v_{\absolutevalue{e}})\}$ included. Finally, we apply the evaporation operation to disconnect the hyperedge from vertex $u$, $\forall u\in V(\alpha(e))\backslash\{\alpha(v_{1}),\alpha(v_{2}),\dots,\alpha(v_{\absolutevalue{e}})\}$. The resulting connected component of vertices $\{\alpha(v)\}_{v\in V(H)}$ is $H^{\prime}$, and it is straightforward to show that $H$ is isomorphic to $H^{\prime}$ with the bijection $f$.
@@
Intuitively, the three conditions in the definition of hypergraph immersion find explanations in parallel to quantum information:
The first condition, concerning vertices—or locality, requires that two vertices in $H$ are distinct if and only if they correspond to different vertices in $G$. As a result, $H$ cannot introduce or merge vertices. This constraint underscores the role of SLOCC, restricting that only qubits associated with different vertices in QN are constrained by (S)LOCC—the free operation that establishes entanglement as a resource [1]. This locality constraint in QN must uphold throughout entanglement routing.
@@
The third condition requires that each hyperedge in $G$ can be utilized only once to derive a hyperedge in $H$. This rule is a direct reflection of the no-cloning theorem [43, *q-no-clone_wz82], asserting that a quantum state cannot be replicated. In other words, once a hyperedge is utilized for entanglement routing, it cannot be reused.
@@
[Uncaptioned image] $K_{3}^{2}$exists a cycle of length $\geq 3$, or two length-$2$ cycles that are edge-disjoint and share exactly one common vertex.
[Uncaptioned image] $K_{3}^{3}$exists a connected component of size $\geq 3$ in $G$.
@@
[Uncaptioned image] $K_{4}^{2}$exists an (ordinary) graph, derived through evaporations on hyperedges in $G$, that is either non-series-parallel [45] or falls into the special cases given in Ref. [46] (see SM).
[Uncaptioned image] $K_{4}^{3}$exists an immersion of one of the $18$ hypergraph variants of $K_{4}^{3}$ in $G$ (see SM).
[Uncaptioned image] $K_{4}^{4}$exists a connected component of size $\geq 4$ in $G$.
@@
Examples.—Let the resulted multipartite QN after entanglement routing be a complete $r$-uniform hypergraph, $H=K_{n}^{r}$, where $n\equiv\left|V(K_{n}^{r})\right|$, and the edge set $E(K_{n}^{r})$ is equal to the set of all size-$r$ subsets in $V(K_{n}^{r})$. In other words, $K_{n}^{r}$ includes all possible hyperedges of size $r$. We present a few necessary and sufficient conditions for $H=K_{n}^{r}$ to be immersed in $G$ in Table 1. We observe that the conditions are straightforward for most small $n$ and $r$, except for $K_{4}^{2}$ and $K_{4}^{3}$, which represent the first nontrivial examples. Indeed, while both $K_{4}^{2}$ and $K_{4}^{3}$ are three-edge-connected, it can be shown that neither one can be immersed in the other [Fig. 2]. This is an explicit example showing that immersion is a partial order, not a total order.
We discuss the criteria for immersion of $K_{4}^{2}$ and $K_{4}^{3}$ further in the SI. These criteria lead to more compelling examples, involving infinite network topologies:
Figure 2: $K_{4}^{2}$ versus $K_{4}^{3}$. 2 Three-edge-connectedness. Both $K_{4}^{2}$ and $K_{4}^{3}$ are three-edge-connected, i.e., every pair of vertices (such as A and B) are linked by at least three edge-disjoint paths (red, blue, orange). But $K_{4}^{2}$ and $K_{4}^{3}$ cannot be immersed into each other. 2 In an infinite honeycomb lattice, there exists an immersion of $K_{4}^{2}$, 2 but not $K_{4}^{3}$. Conversely, in a cycle graph made up of $\geq 4$ vertices and double-edges, while 2 there cannot be an immersion of $K_{4}^{2}$, 2 an immersion of $K_{4}^{3}$ exists.
@@
There is an immersion of $K_{4}^{3}$ but not $K_{4}^{2}$ in a cycle graph consisting of only double-edges and no less than $4$ vertices.
@@
The proof of the existence of an immersion $\alpha$ of $H=K_{4}^{3}$ in a cycle graph consisting of double edges and no less than $4$ vertices ($G$) can be shown by construction. See Fig. 2 for an example in $G$ of size $7$, where the four vertices in $H$ are mapped to $A,B,C,D$ in $G$. The generalisation to $G$ of an arbitrary size $n$ is straightforward.
@@
For $H=K_{4}^{2}$, assume by contradiction that such immersion exists, denoted by $\alpha$. Assume w.l.o.g. the four vertices $A^{\prime},B^{\prime},C^{\prime},D^{\prime}$ in $H$ are mapped to $A,B,C,D$, respectively, that are consecutive anti-clockwise in $G$. The presence of edges $A^{\prime}B^{\prime},A^{\prime}C^{\prime},A^{\prime}D^{\prime}$ implies three edge-disjoint paths between $A,B$, $A,C$, and $A,D$, respectively. Since node $C$ can only reach node $A$ through nodes $B$ or $D$, assume w.l.o.g. that $B\in\alpha(A^{\prime}C^{\prime})$. Hence, there are two edges that are incident on node $B$ in $\alpha(A^{\prime}C^{\prime})$, and there is also one edge that is incident on node $B$ in $\alpha(A^{\prime}B^{\prime})$. Then, there is only one edge in $G$ that is incident on node $B$ but is not in $\alpha(A^{\prime}B^{\prime})$ or $\alpha(A^{\prime}C^{\prime})$. However, $\alpha(B^{\prime}C^{\prime})$ and $\alpha(B^{\prime}D^{\prime})$ imply that there are two edges that are incident on node $B$ but are not in $\alpha(A^{\prime}B^{\prime})$ or $\alpha(A^{\prime}C^{\prime})$, which leads to contradiction. ∎
@@
These examples have profound implications for QN designs. For instance, now we know that the 2D lattice in Fig. 2 is impossible for simultaneous secret sharing among every three parties in any four-party group through entanglement routing, unless additional quantum channels [47] or memories [48, *q-netw-mem_mpnk24] are introduced to the QN to create more links. Similarly, the quasi-1D structure in Fig. 2 also has inherent topological limitations. To conclude, the hypergraph immersion problem we introduce captures the topological landscape of entanglement routing in QNs, which could allow for analyzing very large scales (due to its polynomial complexity [32]).
@@
Discussion.—In classical communication, information transmission is fundamentally directional [50], reflecting the causality between the sender and receiver. Hence, a classical broadcast can only disseminate information from one-to-many, not many-to-one. In contrast, quantum entanglement lacks a causal structure, as suggested by the no-communication theorem [51], which can be used to transmit classical information. This absence of causality underscores the unique interpretation of multipartite entangled states as undirected multiedges, a pure quantum analog. Also, note that a variant of immersion involves labeled vertices in both hypergraphs $G$ and $H$, corresponding to the establishment of multipartite entanglement between specified vertices. In this labeled scenario, identifying an immersion of $H$ in $G$ is simpler than in the unlabeled case. We leave this direction for future exploration.
Appendix A Supplemental Materials
Appendix B 1. Terminology
@@ Appendix A Supplemental Materials
A graph $G$ consists of vertices and edges, where we denote by $V(G)$, $E(G)$ for the set of vertices and edges, respectively. We consider undirected graphs in this paper, where multiple edges may be present among the same set of nodes, but loops are ignored. In an ordinary graph, an edge can only connect two vertices, thus $\forall e\in E(G)$, $e=uv$ with $u,v\in V(G)$.
@@
Connections between vertices are important in a graph. Specifically, a cut point in a connected graph $G$ is a vertex whose removal disconnects $G$. Two vertices in G are said to be biconnected if they cannot be disconnected by the removal of any cut point. A biconnected component of $G$ is the subgraph induced by a maximal set of pairwise biconnected vertices.
@@
Two vertices are said to be three-edge-connected if there are at least three edge-disjoint paths between them. A three-edge-connected component of $G=(V,E)$ is a graph $G^{\prime}=(V^{\prime},E^{\prime})$ where $V^{\prime}\subseteq V$ is a maximal set of vertices that are pairwise three-edge-connected in $G$. $E^{\prime}$ contains all edges induced by $V^{\prime}$ plus a (possibly empty) set of virtual edges defined as follows: for $\{u,v\}\subseteq V^{\prime}$, a virtual edge $uv$ is added to $E^{\prime}$ for each distinct $\{x,y\}\subseteq V-V^{\prime}$ such that $ux$ and $vy$ form a cut edge pair in $G$. Note that, due to the possible preference of virtual edges, a three-edge-connected component will not necessarily be a subgraph.

### v2

# Multipartite Entanglement Routing as a Hypergraph Immersion Problem

@@ Abstract
Multipartite entanglement, linking multiple nodes simultaneously, is a higher-order correlation that offers advantages over pairwise connections in quantum networks (QNs). Creating reliable, large-scale multipartite entanglement requires entanglement routing, a process that combines local, short-distance connections into a long-distance connection, which can be considered as a transformation of network topology. Here, we address the question of whether a QN can be topologically transformed into another via entanglement routing. Our key result is an exact mapping from multipartite entanglement routing to Nash-Williams’s graph immersion problem, extended to hypergraphs. This generalized hypergraph immersion problem introduces a partial order between QN topologies, permitting certain topological transformations while precluding others, offering discerning insights into the design and manipulation of higher-order network topologies in QNs.
@@
Quantum entanglement, a pivotal quantum resource [1], allows for the simultaneous correlation of multiple qubits. Multipartite entanglement offers notable benefits, in terms of both efficiency and scalability, over bipartite entanglement in a variety of quantum information applications such as sharing secrets among multiple parties [2] and reducing the memory requirements for entangling qubits [3]. These benefits make multipartite entanglement a necessary building block for future quantum networks (QN) [4, 5, *multipartite-q-netw_ctpv21, 7]. In a QN, each vertex (node) typically represents a local assembly of qubits entangled with qubits from other vertices, and each edge (link) signifies an entangled state between qubits that belong to different vertices [8, *conpt_mgh21]. In this setting, multipartite entanglement is most effectively represented as a higher-order interaction, or in the language of hypergraph theory, as a hyperedge [10]. Contrasting with ordinary edges that connect just two vertices, hyperedges can connect several vertices at once—a feature that has wider applications in not only physics but also chemistry [11], biology [12], and social sciences [13, *krumov2011collab, *bianconi2021hyper, *Kwang2023core].
@@
In QNs, vertices not directly connected by an edge can still be entangled through a fundamental operation known as entanglement routing [17, *q-netw-route_pkttjbeg19, *q-netw-route_lphnm20, 20], provided that there is a path of edges connecting the vertices. This operation, however, utilizes all the edges along the path, subsequently deleting them from the QN, leading to a dynamic process called path percolation [21]. As a result, each entanglement routing operation uniquely modifies the QN’s topology in an irreversible manner, permitting certain types of topological transformations while precluding others, raising the question: can a QN’s topology be transformed from one to another through a series of entanglement routing operations? It is also unknown how this question extends to hypergraph QNs due to the ambiguous definition of paths involving multipartite entanglement [20]. Understanding this question is key to grasping the complexity of many-body entanglement structures (e.g., matrix product states [22], which can be reformulated as QN representations [7, 23]). Moreover, the insights can be applied to evaluate the effectiveness of qubit layouts (e.g., the hex lattices used by IBM [24]). This is especially relevant as small-scale multipartite-entanglement QNs have been implemented on superconducting [25, *chen2014qubit] and optical [27, *kok2007linear] platforms.
@@
Here, we address this question of entanglement routing, involving both edges and hyperedges, by establishing an exact mapping to a topological problem on hypergraphs. We first identify and simplify entanglement routing into two hypergraph rules by considering the routing operation for multipartite entanglement. Then, we show that the corresponding topological problem can be viewed as a generalization of the well-established graph immersion problem [29], which, to our best knowledge, has not been extended to hypergraphs. This aptly allows us to formalize our problem as hypergraph immersion, which establishes a partial order [30] between hypergraph topologies. Following graph immersion [31], we also prove in an accompanying work that the algorithmic complexity of solving hypergraph immersion is polynomial [32]. This unique bridge between quantum information and graph theory may provide some useful insights into the efficient design of multipartite QNs.
@@
Figure 1: Hypergraph immersion. (a) The “coalescence” operation merges two hyperedges that share at least one vertex, resulting in a new hyperedge that joins all vertices from the original two. (b) The “dewetting” operation detaches a vertex from a hyperedge, reducing the hyperedge’s size by one. (c) The desired final QN topology can be reached from the initial topology through a sequence of operations, provided that the final topology is immersed within the original structure.
@@
These states, albeit varying in size ($r\geq 2$), always have the lowest nontrivial generalized Schmidt rank [33, *eisert2001schmidt] to exhibit entanglement ($\text{rank}=2$). GHZ states are prevalent in quantum systems, such as the random transverse-field quantum Ising model, where the critical ground state consists exclusively of GHZ-state magnetic domains [35]. The Ising ground state naturally facilitates a QN interpretation using complex network theory [36].
How do we entangle multiple vertices by routing via several hyperedges? Given, for example, two GHZ states of $r_{1}$ and $r_{2}$ qubits, as well as at least one common vertex that shares two qubits separately from the two states, a simple projector on the two qubits,
@@
on the vertex, yielding a smaller GHZ state with $r-1$ qubits. With appropriate rotations of |$\ket{0}$ and |$\ket{1}$, these two operations are also applicable to general Schmidt-rank-2 states of the form $|\alpha_{1}\otimes\alpha_{2}\otimes...\otimes\alpha_{n}\rangle+z|\beta_{l}\otimes\beta_{2}\otimes...\otimes\beta_{n}\rangle$ [37]. An example of such is the cat state, commonly used in continuous-variable optical communications [38].
@@
Coalescence and dewetting operations.—The topological emphasis of Eqs. (2) and (3) suggests that we can simplify them as hypergraph operations. Consider two hypergraphs $H$ and $G$, each comprising sets of vertices $V(H)$ and $V(G)$, and sets of hyperedges $E(H)$ and $E(G)$, respectively. Both $H$ and $G$ are loopless  [^1] We note that a loop in (hyper)graphs means an edge contains the same vertex more than once. The case where an edge only includes one vertex is allowed. and undirected, but may include multi-hyperedges. We claim that the feasibility of deriving $H$ from $G$ through entanglement routing is equivalent to asserting that $H$ can be derived from $G$ through the following hypergraph operations:
@@
  • Dewetting: detaching a hyperedge from one vertex that it is incident on.
One can visualize vertices in the hypergraph as spatially distributed objects (pink) that are “hydrophilic” (Fig. 1). Correspondingly, each hyperedge $e$ can be visualized as a water droplet, touching those hydrophilic objects (vertices) that are incident with $e$. In this physical analogy, the coalescence operation results in the merging of two droplets that share contact with a common vertex $v$, forming a larger droplet [Fig. 1]; the dewetting operation simulates the reduction in the volume of a droplet and the loss of its connections to the vertices it was previously in contact with [Fig. 1]. This physical analogy motivates our naming of the two operations.
Note that both operations are irreversible: the coalescence reduces the number of hyperedges by one; the dewetting reduces the number of vertices incident with the hyperedge by one. Hence, for $H$ to be derived from $G$ through these operations, a necessary condition is $\left|V(H)\right|\leq\left|V(G)\right|$ and $\left|E(H)\right|\leq\left|E(G)\right|$, i.e., $H$ must not possess more vertices or hyperedges than $G$. In general, determining whether $H$ can be derived from $G$ poses a nontrivial problem. In the following, we present an alternative definition in parallel to these operations.
@@
In other words, $\alpha$ is an injective mapping from vertices in $H$ to vertices in $G$, and from hyperedges in $H$ to edge-disjoint connected subgraphs in $G$. The formulation of $\alpha$ mirrors the graph immersion as outlined in Ref. [29], which has a similar definition that maps edges in $H$ to edge-disjoint paths in $G$, but with $H$ and $G$ being ordinary graphs. For ordinary edges, the function $\alpha$ reduces to the definition of graph immersion. More (hyper)graph theory terminologies can be found in SM, Section 1.
@@
An immersion $\alpha$ of hypergraph $H$ in hypergraph $G$ exists if and only if $H$ can be obtained from a subgraph of $G$ by a sequence of coalescence and dewetting operations.
@@
If an immersion $\alpha$ exists, we can show that $H$ is isomorphic to a hypergraph $G^{\text{cl;dw}}$ obtained by applying the coalescence and dewetting operations on a subgraph of $G$, where the bijection corresponding to the isomorphism $f:V(H)\to V(H^{\prime})$ takes the same value as $\alpha$, i.e., $\forall v\in V(H)$, $f(v)=\alpha(v)$.
We build $G^{\text{cl;dw}}$ as follows. For each $e\in E(H)$ with distinct ends $v_{1},v_{2},\dots,v_{\left|e\right|}$, we apply the coalescence operation to every adjacent pair of hyperedges in the corresponding connected component $\alpha(e)\subseteq G$, until we obtain one single hyperedge incident to all vertices in $\{\alpha(v_{1}),\alpha(v_{2}),\dots,\alpha(v_{\left|e\right|})\}$. The coalescence operation is possible since $\alpha(e)$ is connected. Then, we apply the dewetting operation to detach the single hyperedge from vertex $u$, $\forall u\in V(\alpha(e))\backslash\{\alpha(v_{1}),\alpha(v_{2}),\dots,\alpha(v_{\left|e\right|})\}$. Doing this for each edge-disjoint $\alpha(e)$, $\forall e\in E(H)$, we denote the resulted hypergraph, a collection of single hyperedges derived from every $\alpha(e)$, by $G^{\text{cl;dw}}$. It is straightforward to show that $H$ is isomorphic to $G^{\text{cl;dw}}$ with the bijection $f$.
@@
Intuitively, the three conditions in the definition of hypergraph immersion can be understood in parallel with concepts in quantum information:
The first condition, concerning vertices—or locality, requires that two vertices in $H$ are distinct if and only if they correspond to different vertices in $G$. In other words, $H$ cannot introduce or merge vertices. This constraint underscores the fact that only qubits associated with different QN vertices are restricted by (S)LOCC—the free operation that establishes entanglement as a resource [1]. This locality constraint in QN must be upheld throughout entanglement routing.
@@
The third condition requires that each hyperedge in $G$ can be utilized only once to derive a hyperedge in $H$. This rule is a direct reflection of the no-cloning theorem [44, *q-no-clone_wz82], asserting that a quantum state cannot be replicated. In other words, once a hyperedge is utilized for entanglement routing, it cannot be reused.
@@
[Uncaptioned image] $K_{3}^{2}$exists a Berge cycle [46] of length $\geq 3$, or two length-$2$ Berge cycles that are edge-disjoint and share exactly one common vertex.
[Uncaptioned image] $K_{3}^{3}$exists a connected subgraph of size $\geq 3$ in $G$.
@@
[Uncaptioned image] $K_{4}^{2}$exists an (ordinary) graph, derived through dewetting only on hyperedges in $G$, that is either non-series-parallel [47] or falls into the special cases given in Ref. [48] (SM, Section 2.1).
[Uncaptioned image] $K_{4}^{3}$exists a “restricted” graph immersion of one of $18$ ordinary graph variants of $K_{4}^{3}$ (SM, Section 2.2).
[Uncaptioned image] $K_{4}^{4}$exists a connected subgraph of size $\geq 4$ in $G$.
@@
Figure 2: $K_{4}^{2}$ versus $K_{4}^{3}$. 2 Both $K_{4}^{2}$ and $K_{4}^{3}$ are three-edge-connected, i.e., every pair of vertices (e.g., A and B) are connected by three edge-disjoint paths (red/pink, blue/navy, yellow/orange). But $K_{4}^{2}$ and $K_{4}^{3}$ cannot be immersed into each other. 2 In an infinite honeycomb lattice, there exists an immersion of $K_{4}^{2}$, 2 but not $K_{4}^{3}$. 2 Conversely, in a cycle graph made up of $\geq 4$ vertices and double-edges, there is no immersion of $K_{4}^{2}$ 2 but $K_{4}^{3}$.
Examples.—Let the resulted multipartite QN after entanglement routing be a complete $r$-uniform hypergraph, $H=K_{n}^{r}$, where $n\equiv\left|V(K_{n}^{r})\right|$, and the edge set $E(K_{n}^{r})$ is equal to the set of all size-$r$ subsets in $V(K_{n}^{r})$. In other words, $K_{n}^{r}$ includes all possible hyperedges of size $r$. We present the first few necessary and sufficient conditions for $H=K_{n}^{r}$ to be immersed in $G$ in Table 1. We observe that the conditions are straightforward for most small $n$ and $r$, except for $K_{4}^{2}$ and $K_{4}^{3}$, which represent the first nontrivial examples. Indeed, while both $K_{4}^{2}$ and $K_{4}^{3}$ are three-edge-connected (SM, Section 1), it can be shown that neither one can be immersed in the other [Fig. 2]. This is an explicit example showing that immersion is a partial order, not a total order.
We discuss the criteria for immersion of $K_{4}^{2}$ and $K_{4}^{3}$ further in the SM. These criteria lead to more compelling examples, involving infinite network topologies:
@@
There is an immersion of $K_{4}^{3}$ but not $K_{4}^{2}$ in a cycle graph consisting of no less than four vertices and only double-edges.
@@
The proof of the existence of an immersion $\alpha$ of $H=K_{4}^{3}$ in a cycle graph consisting of double edges and no less than $4$ vertices ($G$) can be shown by construction. See Fig. 2 for an example in $G$ of size $7$, where the four vertices in $H$ are mapped to $A,B,C,D$ in $G$. The generalization to $G$ of an arbitrary size $n$ is straightforward.
@@
For $H=K_{4}^{2}$, assume by contradiction that such immersion exists, denoted by $\alpha$. Assume w.l.o.g. the four vertices $A^{\prime},B^{\prime},C^{\prime},D^{\prime}$ in $H$ are mapped to $A,B,C,D$, respectively, that are consecutive anti-clockwise in $G$. The presence of edges $A^{\prime}B^{\prime},A^{\prime}C^{\prime},A^{\prime}D^{\prime}$ implies three edge-disjoint paths between $A,B$, $A,C$, and $A,D$, respectively. Since node $C$ can only reach node $A$ through nodes $B$ or $D$, assume w.l.o.g. that $B\in\alpha(A^{\prime}C^{\prime})$. Hence, there are two edges that are incident on node $B$ in $\alpha(A^{\prime}C^{\prime})$, and there is also one edge that is incident on node $B$ in $\alpha(A^{\prime}B^{\prime})$. Then, there is only one edge in $G$ that is incident on node $B$ but is not in $\alpha(A^{\prime}B^{\prime})$ or $\alpha(A^{\prime}C^{\prime})$. However, $\alpha(B^{\prime}C^{\prime})$ and $\alpha(B^{\prime}D^{\prime})$ imply that there are two edges that are incident on node $B$ but are not in $\alpha(A^{\prime}B^{\prime})$ or $\alpha(A^{\prime}C^{\prime})$, which leads to a contradiction. ∎
@@
These examples have profound implications for QN designs. For instance, we have shown that the 2D lattice in Fig. 2 is impossible for simultaneous secret sharing among every three parties in any four-party group through entanglement routing, unless additional quantum channels [49] or memories [50, *q-netw-mem_mpnk24] are introduced to the QN to create more links. Similarly, the quasi-1D structure in Fig. 2 also has inherent topological limitations. To conclude, the hypergraph immersion problem we introduce captures the topological landscape of entanglement routing in QNs, which could allow for analyzing very large scales (due to its polynomial complexity [40]).
@@
Discussion.—In classical communication, information transmission is fundamentally directional [52], reflecting the causality between the sender and receiver. Hence, a classical broadcast can only disseminate information from one-to-many, not many-to-one. In contrast, quantum entanglement lacks a causal structure to transmit classical information, as suggested by the no-communication theorem [53]. This absence of causality underscores the unique interpretation of multipartite entangled states as undirected multiedges, a pure quantum analog. Also, it is worth noting that a variant of immersion involves labeled vertices in both hypergraphs $G$ and $H$, corresponding to the establishment of multipartite entanglement between specified vertices. In this labeled scenario, identifying an immersion of $H$ in $G$ is usually simpler than in the unlabeled case. We leave this direction for future exploration.
Supplemental Materials
Section I Preliminaries
@@
A graph $G$ consists of vertices and edges. We denote by $V(G)$, $E(G)$ for the set of vertices and edges, respectively. We consider undirected graphs in this paper, where multiple edges may be present among the same set of vertices, but loops, i.e., an edge starting and ending at the same node, are ignored. In an ordinary graph, an edge can only connect two vertices, thus $\forall e\in E(G)$, $e=uv$ with $u,v\in V(G)$.
@@
End-to-end connections between vertices are important in a graph. Specifically, a cut point in a connected graph $G$ is a vertex whose removal disconnects $G$. Two vertices in $G$ are said to be biconnected if they cannot be disconnected by the removal of any cut point. A biconnected component of $G$ is the subgraph induced by a maximal set of pairwise biconnected vertices.
@@
Two vertices are said to be three-edge-connected if there are at least three edge-disjoint paths between them. A three-edge-connected component of $G=(V,E)$ is a graph $G^{\prime}=(V^{\prime},E^{\prime})$ where $V^{\prime}\subseteq V$ is a maximal set of vertices that are pairwise three-edge-connected in $G$, and $E^{\prime}$ contains all edges induced by $V^{\prime}$ plus a (possibly empty) set of virtual edges defined as follows: for $\{u,v\}\subseteq V^{\prime}$, a virtual edge $\{u,v\}$ is added to $E^{\prime}$ whenever there exist a pair of distinct nodes $\{x,y\}$ that are not in $V^{\prime}$, such that removing edges $\{u,x\}$ and $\{v,y\}$ disconnects the graph. The three-edge-connected component better represents the connectivity of the nodes, projected onto a smaller graph. Note that, due to the possible presence of virtual edges, a three-edge-connected component will not necessarily be a subgraph.